# Generalized $T\bar{T}$-like flows for scalar theories in two dimensions

### v1

@@ 4 General root $T\bar{T}$-like flows in two dimensions
At the $\lambda^{4}$ order, the differential equation arising from the root deformation 4.7 is quite lengthy. We solve this equation for the specific case where $e_{1}=-\frac{1}{4}$, and determine the coefficient $d_{16}$ as follows:
${d}_{16}$ $=$ $(-4n_{2}^{2}-n_{1}n_{3})\bigl{(}\cosh(\gamma)\bigr{)}^{3}\bigl{(}5\cosh(2\gamma)-3\bigr{)}+\tfrac{1}{8}n_{1}^{4}\cosh(\gamma)\bigl{(}-29+100\cosh(2\gamma)+65\cosh(4\gamma)\bigr{)}$ (4.13)
@@
In this section, we employed the perturbation approach to examine the general form of two-dimensional scalar theories up to order $\lambda^{4}$. We expressed this as the Lagrangian in 4.4 and identified constraints on the coefficients $d_{n}$ to satisfy the $SO(N)\times SO(N)$ duality invariant condition in the differential form 4.2. This process led us to derive the Lagrangian in 4.5. Additionally, we determined the $\gamma$ dependence of the unfixed coefficients $d_{n}$ by imposing the constraint that the theory adheres to a general root flow equation of the form 4.7. By substituting these coefficients 4.9, 4.11, 4.12, and 4.13 into the Lagrangian in 4.5, we obtained the most general form of the Lagrangian with two coupling constants up to order $\lambda^{4}$, ensuring that it is both $SO(N)\times SO(N)$ duality invariant and consistent with the root flow equation.
@@ 5 Generalized irrelevant flow equation and commutability
Given the coefficients $e_{n}$ as $e_{0}=\frac{1}{2}$, $e_{1}=-\frac{1}{4}$, and $e_{2}=e_{3}=...=e_{n}=0$, the root flow equation in 4.7 will take the standard form of 1.5 in two dimensions. Now, with this choice for $e_{n}$ and the substitution of the coefficient $d_{n}$ in the general Lagrangian 4.5, we obtain a general Lagrangian that satisfies the root flow equation $\frac{\partial\mathcal{L}(\lambda,\gamma)}{\partial\gamma}=\mathcal{R}_{\gamma}$. This Lagrangian is as follows:
@@
$+\lambda^{3}\Bigl{(}-2n_{1}^{3}(-P_{1}+\mathcal{P})(P_{1}+\mathcal{P})\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}+\tfrac{1}{4}n_{3}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{4}$
@@
$+\tfrac{1}{4}n_{1}^{4}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}\bigl{(}29P_{1}(-P_{1}+\mathcal{P})(P_{1}+\mathcal{P})\cosh(\gamma)+13P_{1}(P_{1}^{2}+3\mathcal{P}^{2})\cosh(3\gamma)$
@@
We can explicitly verify that the Lagrangian 5 possesses a root flow equation. With respect to root deformation, the general Lagrangian in 5 simplifies to the two-dimensional ModMax Lagrangian when $\lambda=0$. This general Lagrangian includes an unfixed coefficient $n_{i}$ at each order of $\lambda$. Specifically, at the order $\lambda^{4}$, it contains four unfixed coefficients: $n_{1}$, $n_{2}$, $n_{3}$, and $n_{4}$. We can derive theories that exhibit a root flow equation by determining these constant coefficients. For instance, by setting the constants $n_{i}$ to $n_{1}=\frac{1}{8}$, $n_{2}=0$, $n_{3}=\frac{5}{256}$, and $n_{4}=0$, we can expand the general two-dimensional Generalized Scalar ModMax (GSMM) theory in Eq. 1.7 up to the order $\lambda^{4}$.
This section will discuss the importance of identifying the general irrelevant flow equation for Lagrangian 5 that commutes with the root deformation in 1.5. By taking the derivative of Lagrangian 5 with respect to $\lambda$ and comparing it with structures A.6 and A.7, we can derive the flow equation of the Lagrangian 5 using a perturbation approach up to the $\lambda^{4}$ order as follows:
@@
$C_{0}=n_{1},\,\,\,\,\,C_{1}=\frac{64}{3}n_{2},\,\,\,\,\,C_{2}=-\frac{17}{16}-\frac{32768}{9}n_{2}^{2}+48\,n_{3},$ (5.3)
@@
The irrelevant flow equation 5.2 is a general flow equation applicable to all theories with arbitrary $n_{i}$. The coefficients $C_{m}$ of this equation explicitly depend on the coefficients of $n_{i}$. We explicitly derive the irrelevant $T\bar{T}$-like deformation of the Lagrangian 5 in 5.2, which represents the general irrelevant $T\bar{T}$-like deformation. This deformation commutes with the root deformation in 1.5 as follows:
@@
In this paper, we shown that operators ${\cal O}_{\lambda}=\sum_{m=0}^{\infty}C_{m}Y^{\tfrac{m}{2}}X^{1-\tfrac{m}{2}}$ and ${\cal R}_{\gamma}$ exhibit commutativity. Consequently, it can be explicitly verified that the double-flow equation: $\partial_{\lambda}\partial_{\gamma}\mathcal{L}(\lambda,\gamma)$, holds for the Lagrangian 5 up to the order of $\lambda^{4}$ in the identity of $\partial_{\lambda}\partial_{\gamma}\mathcal{L}(\lambda,\gamma)-\partial_{\gamma}\partial_{\lambda}\mathcal{L}(\lambda,\gamma)=0$. This property is illustrated in Fig. 2.
@@
The Lagrangian of 5 can be categorized into two classes, each characterized by irrelevant $T\bar{T}$-like deformations. We can consider $n_{i}$ as $i=even$ and $i=odd$. The first class includes $n_{i}$ where $n_{even}=0$ and $n_{odd}\neq 0$ These theories exhibit irrelevant flow equations with integer powers of the energy-momentum tensor. The second class consists of theories with $n_{even}\neq 0$ and $n_{odd}\neq 0$, featuring flow equations with fractional powers of the energy-momentum tensor as 5.2. Interestingly, the first class is a special subgroup of the second class. In other words, when we study theories involving $n_{even}=0$ and $n_{odd}\neq 0$, it automatically transforms into $C_{odd}=0$, leaving only the integer powers of the two structures, $X$ and $Y$.
6 Conclusions and perspects
@@
Furthermore, we identify a general form for these flow equations that commute with each other. This non-trivial commutativity implies that the order in which the deformations are applied does not affect the outcome, which is essential for maintaining duality invariance. The commutative nature of flow equations underpins the robustness of the theoretical framework and opens new avenues for further research in this area of theoretical physics. Potential applications of these results extend to theories of chiral $p-$forms [35, 47, 64, 65], supersymmetry [66, 67], and gravitational theories involving the Ricci flow equations [68, 69]. In [69], it is briefly discussed that the general marginal deformation 4.7 at the Lagrangian level corresponds to a ”pure change of metric.”
@@
The generalized theory unifies different deformation paths governed by two coupled constants, $\lambda$ and $\gamma$, each controlling various aspects of the deformation structure. Importantly, our results reveal the rich underlying nature of these models. Future research could explore higher-order solutions and their implications for integrability, quantum consistency, and potential connections to string theory via compactification mechanisms. In addition, it is intriguing to identify an operator derived from the energy-momentum tensor that characterizes double-flow equation. We also want to discover closed forms for the general Lagrangians discussed in this paper.
@@ Acknowledgments
The authors are grateful to Dmitri Sorokin and Roberto Tateo for valuable discussions. S.H. acknowledges financial support from Ningbo University, the Max Planck Partner Group, and the Natural Science Foundation of China Grants No. 12475053, No. 12075101, No. 12235016, No. 12347209. H.O. is supported by the National Natural Science Foundation of China, Grant No. 12205115, and by the Science and Technology Development Plan Project of Jilin Province of China, Grant No. 20240101326JC.
@@ Appendix A Details of the Lagrangian, and the Energy-Momentum Tensor
This appendix provides a detailed account of some of the perturbation calculations from the paper, which are extensive. Given that general Lagrangian 3.1 is valid in the PDE 1.10, there exists a relationship between the coefficients $a_{n}$ at each order of $\lambda$, resulting in the following correlation between the coefficients $a_{n}$ as:
$\Big{(}a_{1}\to-\tfrac{1}{2}\Big{)};\,\,\,\,\,\,\,\,\Big{(}a_{3}\to-2a_{2}\Big{)};\,\,\,\,\,\,\,\,\Big{(}a_{4}\to 4{a_{2}}^{2},a_{5}\to-8{a_{2}}^{2}\Big{)};\,\,\,\,\,\,\,\,\Big{(}a_{6}{}\to\tfrac{1}{4}\bigl{(}32{a_{2}}^{3}-a_{7}{}\bigr{)}\,\,,\,\,a_{8}{}\to-32{a_{2}}^{3}-a_{7}{}\Big{)};$
@@
$a_{17}\to 2\bigl{(}-29056{a_{2}}^{6}-1376{a_{2}}^{3}a_{7}{}-3{a_{7}}^{2}\ +12a_{2}{}a_{13}{}\bigr{)}\,\,,\,\,a_{18}\to-4\bigl{(}-35712{a_{2}}^{6}-1696{a_{2}}^{3}a_{7}{}-3{a_{7}}^{2}+12a_{2}{}a_{13}{}\bigr{)}$
@@
[…]
$a_{29}\to 32\bigl{(}-2114944{a_{2}}^{8}-105920{a_{2}}^{5}a_{7}{}-484{a_{2}}^{2}{a_{7}}^{2}+688{a_{2}}^{3}a_{13}{}+a_{7}{}a_{13}{}-2a_{2}{}a_{21}{}\bigr{)}\Big{)}.$ (A.1)
The details of the energy-momentum tensor are provided in Eq. 3.3, involving two functions, ${\cal F}(P_{1},P_{2},\lambda)$ and ${\cal G}(P_{1},P_{2},\lambda)$, as follows:
@@
$+2\lambda^{7}P_{1}{}\Bigl{(}a_{21}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}-56a_{2}{}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}+112{a_{2}}^{2}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}$
$-3584{a_{2}}^{4}a_{7}{}\bigl{(}3{P_{1}}^{6}-46{P_{1}}^{2}{P_{2}}^{2}+68{P_{2}}^{3}\bigr{)}-32768{a_{2}}^{7}\bigl{(}7{P_{1}}^{6}-103{P_{1}}^{2}{P_{2}}^{2}+152{P_{2}}^{3}\bigr{)}\Bigr{)}+{\cal O}(\lambda^{8}).$
@@
${\cal G}(P_{1},P_{2},\lambda)$ $=$ $8\lambda a_{2}{}+32\lambda^{2}{a_{2}}^{2}P_{1}{}+\lambda^{3}\Bigl{(}-4a_{7}{}{P_{1}}^{2}+8\bigl{(}32{a_{2}}^{3}+a_{7}{}\bigr{)}P_{2}{}\Bigr{)}$
@@
$+32{a_{2}}^{3}a_{7}{}\bigl{(}43{P_{1}}^{4}-212{P_{1}}^{2}P_{2}{}+252{P_{2}}^{2}\bigr{)}+128{a_{2}}^{6}\bigl{(}227{P_{1}}^{4}-1116{P_{1}}^{2}P_{2}{}+1356{P_{2}}^{2}\bigr{)}\Bigr{)}$
@@
$+224{a_{2}}^{3}a_{7}{}\bigl{(}23{P_{1}}^{4}-102{P_{1}}^{2}P_{2}{}+112{P_{2}}^{2}\bigr{)}+1024{a_{2}}^{6}\bigl{(}103{P_{1}}^{4}-456{P_{1}}^{2}P_{2}{}+504{P_{2}}^{2}\bigr{)}\Bigr{)}\biggr{)}+{\cal O}(\lambda^{8}).$
The Eq. 4.6 provides the details of the energy-momentum tensor, expressed through the functions ${\cal H}$ and ${\cal U}$ as follows:
@@
×$\times\bigl{(}(-30P_{1}^{2}+38\mathcal{P}^{2})\cosh(\gamma)+(39P_{1}^{2}-5\mathcal{P}^{2})\cosh(3\gamma)+3(-3P_{1}^{2}+5\mathcal{P}^{2})\cosh(5\gamma)$
@@
${\cal U}$ $=$ $\frac{2\sinh(\gamma)}{\mathcal{P}}-\lambda\frac{4d_{4}\bigl{(}\coth(\gamma)\mathcal{P}+P_{1}\bigr{)}}{\mathcal{P}}$
@@
$+d_{4}^{3}\biggl{(}-8\coth(\gamma)\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{4}\mathcal{P}^{2}+\frac{8P_{1}^{2}\mathrm{sech}(\gamma)\Bigl{(}\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{3}\mathcal{P}+2\mathrm{csch}(\gamma)\mathcal{P}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{2}+2P_{1}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{3}\Bigr{)}}{\mathcal{P}}\biggr{)}\Biggr{)}$

### v2

@@ 4 General root $T\bar{T}$-like flows in two dimensions
At the $\lambda^{4}$ order, the differential equation arising from the root deformation (4.7) is quite lengthy. We solve this equation for the specific case where $e_{1}=-\frac{1}{4}$, and determine the coefficient $d_{16}$ as follows:
${d}_{16}$ $=$ $(-4n_{2}^{2}-n_{1}n_{3})\bigl{(}\cosh(\gamma)\bigr{)}^{3}\bigl{(}5\cosh(2\gamma)-3\bigr{)}$
$+\tfrac{1}{8}n_{1}^{4}\cosh(\gamma)\bigl{(}-29+100\cosh(2\gamma)+65\cosh(4\gamma)\bigr{)}$
@@
In this section, we employed the perturbation approach to examine the general form of two-dimensional scalar theories up to order $\lambda^{4}$. We expressed this as the Lagrangian in (4.4) and identified constraints on the coefficients $d_{n}$ to satisfy the $SO(N)\times SO(N)$ duality invariant condition in the differential form (4.2). This process led us to derive the Lagrangian in (4). Additionally, we determined the $\gamma$ dependence of the unfixed coefficients $d_{n}$ by imposing the constraint that the theory adheres to a general root flow equation of the form (4.7). By substituting these coefficients (4.9), (4.11), (4), and (4) into the Lagrangian in (4), we obtained the most general form of the Lagrangian with two coupling constants up to order $\lambda^{4}$, ensuring that it is both $SO(N)\times SO(N)$ duality invariant and consistent with the root flow equation.
@@ 5 Generalized irrelevant flow equation and commutability
Given the coefficients $e_{n}$ as $e_{0}=\frac{1}{2}$, $e_{1}=-\frac{1}{4}$, and $e_{2}=e_{3}=...=e_{n}=0$, the root flow equation in (4.7) will take the standard form of (1.5) in two dimensions. Now, with this choice for $e_{n}$ and the substitution of the coefficient $d_{n}$ in the general Lagrangian (4), we obtain a general Lagrangian that satisfies the root flow equation $\frac{\partial\mathcal{L}(\lambda,\gamma)}{\partial\gamma}=\mathcal{R}_{\gamma}$. This Lagrangian is as follows:
@@
$+\lambda^{3}\Bigl{(}-2n_{1}^{3}(-P_{1}+\mathcal{P})(P_{1}+\mathcal{P})\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}$
$+\tfrac{1}{4}n_{3}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{4}$
@@
$+\tfrac{1}{4}n_{1}^{4}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}$
$\times\bigl{(}29P_{1}(-P_{1}+\mathcal{P})(P_{1}+\mathcal{P})\cosh(\gamma)+13P_{1}(P_{1}^{2}+3\mathcal{P}^{2})\cosh(3\gamma)$
@@
We can explicitly verify that the Lagrangian (5) possesses a root flow equation. With respect to root deformation, the general Lagrangian in (5) simplifies to the two-dimensional ModMax Lagrangian when $\lambda=0$. This general Lagrangian includes an unfixed coefficient $n_{i}$ on each order of $\lambda$. Specifically, on the order $\lambda^{4}$, it contains four unfixed coefficients: $n_{1}$, $n_{2}$, $n_{3}$, and $n_{4}$. We can derive theories that exhibit a root flow equation by determining these constant coefficients. For example, setting the constants $n_{i}$ to $n_{1}=\frac{1}{8}$, $n_{2}=0$, $n_{3}=\frac{5}{256}$, and $n_{4}=0$, we can expand the general two-dimensional Generalized Scalar ModMax (GSMM) theory in Eq. (1.7) up to the order $\lambda^{4}$.
The general integrable theory derived in Eq. (5) contains an undetermined constant at each order of the $\lambda$-expansion, where these coefficients are theory-dependent. As established in the literature, this framework enables the dimensional reduction of four-dimensional electrodynamic theories to two dimensions, with Born-Infeld theory serving as a prime example. Crucially, this reduction preserves causality when applied to four-dimensional theories that respect the causality principle, producing novel two-dimensional integrable systems (see [53, 54] for representative cases). The resulting two-dimensional theories maintain closed-form expressions, and their $\lambda$- expansions are systematically captured by our general Lagrangian in Eq. (5). Specific theories emerge from this unified framework through the appropriate choices of the expansion coefficients. This construction represents a significant advancement as it provides: 1: a perturbative classification of integrable theories with two field variables ($P_{1}$, $P_{2}$), 2: consistent incorporation of two coupling constants ($\gamma$, $\lambda$), and a solid foundation for future investigations in this field.
This section will discuss the importance of identifying the general irrelevant flow equation for Lagrangian (5) that commutes with the root deformation in (1.5). By taking the derivative of Lagrangian (5) with respect to $\lambda$ and comparing it with structures (A) and (A.7), we can derive the flow equation of the Lagrangian (5) using a perturbation approach up to the $\lambda^{4}$ order as follows:
@@
$C_{0}=n_{1},\,\,\,\,\,C_{1}=\frac{64}{3}n_{2},\,\,\,\,\,C_{2}=-\frac{17}{16}-\frac{32768}{9}n_{2}^{2}+48\,n_{3},$
@@
The irrelevant flow equation (5.2) is a general flow equation applicable to all theories with arbitrary $n_{i}$. The coefficients $C_{m}$ of this equation explicitly depend on the coefficients of $n_{i}$. We explicitly derive the irrelevant $T\bar{T}$-like deformation of the Lagrangian (5) in (5.2), which represents the general irrelevant $T\bar{T}$-like deformation. This deformation commutes with the root deformation in (1.5) as follows:
@@
In this paper, we shown that operators ${\cal O}_{\lambda}=\sum_{m=0}^{\infty}C_{m}Y^{\tfrac{m}{2}}X^{1-\tfrac{m}{2}}$ and ${\cal R}_{\gamma}$ exhibit commutativity. Consequently, it can be explicitly verified that the double-flow equation: $\partial_{\lambda}\partial_{\gamma}\mathcal{L}(\lambda,\gamma)$, holds for the Lagrangian (5) up to the order of $\lambda^{4}$ in the identity of $\partial_{\lambda}\partial_{\gamma}\mathcal{L}(\lambda,\gamma)-\partial_{\gamma}\partial_{\lambda}\mathcal{L}(\lambda,\gamma)=0$. This property is illustrated in Fig. (2).
@@
The Lagrangian of (5) can be categorized into two classes, each characterized by irrelevant $T\bar{T}$-like deformations. We can consider $n_{i}$ as $i=even$ and $i=odd$. The first class includes $n_{i}$ where $n_{even}=0$ and $n_{odd}\neq 0$ . These theories exhibit irrelevant flow equations with integer powers of the energy-momentum tensor. The second class consists of theories with $n_{even}\neq 0$ and $n_{odd}\neq 0$, featuring flow equations with fractional powers of the energy-momentum tensor as (5.2). Interestingly, the first class is a special subgroup of the second class. In other words, when we study theories involving $n_{even}=0$ and $n_{odd}\neq 0$, it automatically transforms into $C_{odd}=0$, leaving only the integer powers of the two structures, $X$ and $Y$.
6 Conclusions and perspectives
@@
The resulting theory in (5) exhibits three fundamental properties: first, it contains theory-dependent coefficients at each order of the $\lambda$-expansion; second, it proposes a perturbative framework for classifying integrable systems with two Lorentz-invariant variables ($P_{1}$, $P_{2}$) for arbitrary couplings ($\lambda$, $\gamma$); and third, it unifies known and novel integrable theories under a single consistent formalism. This work thereby outlines a constructive method and suggests a perturbative classification scheme for this class of theories.
Furthermore, we identify a general form for these flow equations that commute with each other. This non-trivial commutativity implies that the order in which the deformations are applied does not affect the outcome, which is essential for maintaining duality invariance. The commutative nature of flow equations underpins the robustness of the theoretical framework and opens new avenues for further research in this area of theoretical physics. Potential applications of these results extend to theories of chiral $p$-forms [37, 52, 69, 70, 71], supersymmetry [72, 73], and gravitational theories involving the Ricci flow equations [74, 75]. In [75], it is briefly discussed that the general marginal deformation  (4.7) at the Lagrangian level corresponds to a ”pure change of metric.”
@@
The generalized theory unifies different deformation paths governed by two coupled constants, $\lambda$ and $\gamma$, each controlling various aspects of the deformation structure. Importantly, our results reveal the rich underlying nature of these models. Future research could explore higher-order solutions and their implications for integrability, quantum consistency, and potential connections to string theory via compactification mechanisms. In addition, it is intriguing to identify an operator derived from the energy-momentum tensor that characterizes double-flow equation. We also aim to derive closed-form expressions for the general Lagrangians discussed in this paper in future work. One approach to obtaining closed-form Lagrangians involves the Courant-Hilbert method. The PDE (1.10) reduces, through appropriate changes of variables, to the Courant-Hilbert equation. The most general solution to this equation is given by the Courant-Hilbert function $\ell(\tau)$, which depends on a parameterized real variable $\tau$. An alternative perspective on this equivalence can be obtained using the auxiliary field formalism introduced in [13].
@@ Acknowledgments
The authors are grateful to Dmitri Sorokin, Roberto Tateo and Karapet Mkrtchyan for valuable discussions. S.H. acknowledges financial support from Ningbo University, the Max Planck Partner Group, and the Natural Science Foundation of China Grants No. 12475053, No. 12075101, No. 12235016, No. 12347209. H.O. is supported by the National Natural Science Foundation of China, Grant No. 12205115, and by the Science and Technology Development Plan Project of Jilin Province of China, Grant No. 20240101326JC.
@@ Appendix A Details of the Lagrangian, and the Energy-Momentum Tensor
This appendix provides a detailed account of some of the perturbation calculations from the paper, which are extensive. Given that general Lagrangian (3.1) is valid in the PDE (1.10), there exists a relationship between the coefficients $a_{n}$ at each order of $\lambda$, resulting in the following correlation between the coefficients $a_{n}$ as:
$\Big{(}a_{1}\to-\tfrac{1}{2}\Big{)};\,\,\,\,\,\,\,\,\Big{(}a_{3}\to-2a_{2}\Big{)};\,\,\,\,\,\,\,\,\Big{(}a_{4}\to 4{a_{2}}^{2},a_{5}\to-8{a_{2}}^{2}\Big{)};\,\,\,\,\,\,\,\,$ (A.1)
$\Big{(}a_{6}{}\to\tfrac{1}{4}\bigl{(}32{a_{2}}^{3}-a_{7}{}\bigr{)}\,\,,\,\,a_{8}{}\to-32{a_{2}}^{3}-a_{7}{}\Big{)};$
@@
$a_{17}\to 2\bigl{(}-29056{a_{2}}^{6}-1376{a_{2}}^{3}a_{7}{}-3{a_{7}}^{2}\ +12a_{2}{}a_{13}{}\bigr{)}\,\,$,\,\,
$a_{18}\to-4\bigl{(}-35712{a_{2}}^{6}-1696{a_{2}}^{3}a_{7}{}-3{a_{7}}^{2}+12a_{2}{}a_{13}{}\bigr{)}$
@@
$a_{28}\to-\tfrac{64}{3}\bigl{(}-5515136{a_{2}}^{8}-276544{a_{2}}^{5}\ a_{7}{}-1284{a_{2}}^{2}{a_{7}}^{2}+1840{a_{2}}^{3}a_{13}{}+\ 3a_{7}{}a_{13}{}-6a_{2}{}a_{21}{}\bigr{)}$
$a_{29}\to 32\bigl{(}-2114944{a_{2}}^{8}-105920{a_{2}}^{5}a_{7}{}-484{a_{2}}^{2}{a_{7}}^{2}+688{a_{2}}^{3}a_{13}{}+a_{7}{}a_{13}{}-2a_{2}{}a_{21}{}\bigr{)}\Big{)}.$
The details of the energy-momentum tensor are provided in Eq. (3.3), involving two functions, ${\cal F}(P_{1},P_{2},\lambda)$ and ${\cal G}(P_{1},P_{2},\lambda)$, as follows:
@@
$+2\lambda^{7}P_{1}{}\Bigl{(}a_{21}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}-56a_{2}{}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}$
$+112{a_{2}}^{2}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}-3584{a_{2}}^{4}a_{7}{}\bigl{(}3{P_{1}}^{6}-46{P_{1}}^{2}{P_{2}}^{2}+68{P_{2}}^{3}\bigr{)}$
$-32768{a_{2}}^{7}\bigl{(}7{P_{1}}^{6}-103{P_{1}}^{2}{P_{2}}^{2}+152{P_{2}}^{3}\bigr{)}\Bigr{)}+{\cal O}(\lambda^{8}).$ (A.2)
@@
${\cal G}(P_{1},P_{2},\lambda)=8\lambda a_{2}{}+32\lambda^{2}{a_{2}}^{2}P_{1}{}+\lambda^{3}\Bigl{(}-4a_{7}{}{P_{1}}^{2}+8\bigl{(}32{a_{2}}^{3}+a_{7}{}\bigr{)}P_{2}{}\Bigr{)}$ (A.3)
@@
$+32{a_{2}}^{3}a_{7}{}\bigl{(}43{P_{1}}^{4}-212{P_{1}}^{2}P_{2}{}+252{P_{2}}^{2}\bigr{)}$
$+128{a_{2}}^{6}\bigl{(}227{P_{1}}^{4}-1116{P_{1}}^{2}P_{2}{}+1356{P_{2}}^{2}\bigr{)}\Bigr{)}$
@@
$+224{a_{2}}^{3}a_{7}{}\bigl{(}23{P_{1}}^{4}-102{P_{1}}^{2}P_{2}{}+112{P_{2}}^{2}\bigr{)}$
$+1024{a_{2}}^{6}\bigl{(}103{P_{1}}^{4}-456{P_{1}}^{2}P_{2}{}+504{P_{2}}^{2}\bigr{)}\Bigr{)}\biggr{)}+{\cal O}(\lambda^{8}).$
[…]
@@
$\times\bigl{(}(-30P_{1}^{2}+38\mathcal{P}^{2})\cosh(\gamma)+(39P_{1}^{2}-5\mathcal{P}^{2})\cosh(3\gamma)+3(-3P_{1}^{2}+5\mathcal{P}^{2})\cosh(5\gamma)$
@@
${\cal U}=\frac{2\sinh(\gamma)}{\mathcal{P}}-\lambda\frac{4d_{4}\bigl{(}\coth(\gamma)\mathcal{P}+P_{1}\bigr{)}}{\mathcal{P}}$ (A.5)
@@
$+d_{4}^{3}\biggl{(}-8\coth(\gamma)\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{4}\mathcal{P}^{2}$
$+\frac{8P_{1}^{2}\mathrm{sech}(\gamma)\Bigl{(}\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{3}\mathcal{P}+2\mathrm{csch}(\gamma)\mathcal{P}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{2}+2P_{1}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{3}\Bigr{)}}{\mathcal{P}}\biggr{)}\Biggr{)}$